# Particles before symmetry

### v4

@@ 6 Conclusions
The VB-POV thereby excludes many symmetry-first models. [^39] Most of which are never considered in practice anyway. Indeed, as discussed in Section 5, there is a widespread tacit assumption that the gauge groups of principal bundles simply are the automorphism groups of the fundamental vector spaces. But this assumption is both unwarranted and inconsistently applied. Fortunately, that narrowing of theoretical scope is (currently) a virtue rather than a liability, for our best (current) physics—the Standard Model—falls squarely within its bounds.
To close, the methodological lesson of eliminating the slack between symmetry and geometry yields two broader morals. First, that future developments of gauge theory might do well to begin with structured, fundamental vector bundles and the tensors they carry. Second, that Occam’s razor—if it has an edge here—trims our theories down with the VB-POV.
@@ Appendix A Principal and associated fibre bundles
Given a vector bundle $(E,M,V)$ a covariant derivative $\mathrm{D}$} is an operator:
${\mathrm{D}}:\Gamma(E)\rightarrow\Gamma(T^{*}M\otimes E)$ (A.7)
@@
${\mathrm{D}}(f\kappa)={\mathrm{d}}f\otimes\kappa+f{\mathrm{D}}\kappa$ (A.8)
@@
Let $\mathrm{D}$} be a covariant derivative on $(E,M,V)$, $\mathbf{v}\in E_{x}$ and $\gamma(t)$ a curve in $M$ such that $\gamma(0)=x$. Then we define the parallel transport along $\gamma$ as the unique section $\mathbf{v}_{h}(t)$ of $E|_{\gamma}$ such that:
${\mathrm{D}}_{\gamma^{\prime}{}}\mathbf{v}_{h}=0.$ (A.9)
The existence and uniqueness of this map is guaranteed for $\gamma\subset U$ some open subset of $M$, and it follows from properties of solutions of ordinary differential equations (cf. \@BBOPcitep\@BAP\@BBN(Kobayashi & Nomizu, 1963, Ch. II.2)\@BBCP).
Here $\mathrm{D}$} is an operator, not a tensor. But by introducing a coordinate frame or basis, we can represent it as such. This is the same as for spacetime covariant derivatives, $\nabla$: it is only upon the introduction of a frame or basis that we find an explicit representation.
To define $\Omega$ in terms of $\mathrm{D}$}, we proceed in the usual way:
@@
Given a covariant derivative $\mathrm{D}$} on a vector bundle $E$, the curvature tensor is the unique multilinear bundle map
@@
$\Omega(X,Y)\kappa=\left({\mathrm{D}}_{X}{\mathrm{D}}_{Y}-{\mathrm{D}}_{Y}{\mathrm{D}}_{X}-{\mathrm{D}}_{[X,Y]}\right)\kappa,$
@@
We can see the curvature then as an element of $\Omega:TM\otimes TM\otimes\mathsf{End}(E)$, i.e. as a map valued on the endomorphisms of $E$ (the fiber-linear transformations that are not necessarily automorphisms).
The trace operation is defined as $\mathsf{Tr}:\mathsf{End}(E)\rightarrow C^{\infty}(M)$, and so can be included in a Lagrangian specifying the dynamics of $\nabla$. Since $\mathsf{End}(E)$ is closed under composisiton, we can obtain a Lagrangian 4-form for the action:
@@
It will prove useful to know that, given any vector bundle $(E,M,V)$ the bundle of frames for $E$, called $L(E)$, is itself a principal fibre bundle $(L(E),M,GL(V))$: here elements of $\pi^{-1}(x)$ are linear frames of $E_{x}$, and $G\simeq GL(V)$ acts via $\rho$ on the typical fibres. By construction, $E\simeq L(E)\times_{\rho}V$. Now, for $G^{\prime}{}\subset G\simeq GL(V)$ we can partition the points of each orbit in $P$, $\mathcal{O}_{p}:=Gp$, into orbits of $G^{\prime}{}$. Each such choice gives a principal bundle with group $G^{\prime}{}$ and it induces further structure on the associated vector bundle, e.g. an inner product, by selecting which frames are considered orthonormal. This is also a principal fibre bundle, $(L^{\prime}{}(E),M,G^{\prime}{})$, whose structure group is a proper subgroup of the general linear group, $G^{\prime}{}\subset GL(V)$, taken to be the group that preserves the structure of $V$. This is called the bundle of admissible frames, e.g. of orthonormal frames. Conversely, if $V$ has more than just the structure of a linear vector space, e.g. if it is endowed with an inner product, it will induce a subgroup $G^{\prime}{}\subset GL(V)$ on $P$ that respects that structure.

### v5

@@ 6.2 Two defences
The geometry-first picture has comparative advantages. The symmetry-free explanations just surveyed—of the Higgs mechanism, the Yukawa couplings, and charge quantisation—share a common source: the geometry-first picture eliminates the slack between symmetry and geometry that the principal-bundle picture tolerates. It makes explicit an assumption—that gauge groups coincide with automorphism groups of fundamental bundles—that the principal-bundle picture leaves tacit, and thereby excludes many symmetry-first models—most of which are never considered in practice anyway. [^38] As discussed in Section 5, there is a widespread tacit assumption that gauge groups simply are automorphism groups of fundamental vector spaces. But this assumption is both unwarranted and inconsistently applied.
This is a defence of the geometry-first picture in absolute terms: I judge it to provide superior explanations—a cleaner ontology—for the class of theories to which it applies. The principal gauge field is not altogether denied reality; in the geometry-first picture, the same physical content is carried by the covariant derivative on a specific vector bundle. But I acknowledge that my preference is staked on ontological commitments that not everyone shares.
Of course, the geometry-first picture also has comparative disadvantages. It cannot accommodate gauge theories built from exceptional Lie groups such as $E_{8}$ without losing its explanatory force (Section 5.2). Moreover, the principal-bundle picture’s greater flexibility is a genuine virtue when one wants a single framework covering as wide a class of gauge theories as possible, not just the narrower class that includes the Standard Model.
So let me offer a second defence that does not require my aforementioned commitments. What we can say unconditionally is that, for the Standard Model, the predictions stay the same. Why bother, then, with a reformulation that changes no predictions? As Feynman observed, and as \@BBOPcitet\@BAP\@BBNHunt (2025)\@BBCP has argued in detail, equivalent formulations are not always equally fertile. A genuinely different explanation of the same phenomena is already an epistemic gain—it widens the base from which we move into the unknown. The comparative advantages noted above are sufficient to warrant the geometry-first picture’s consideration as an alternative, even if it does not supersede the principal-bundle picture for all purposes.
Two morals, then. First, future work in gauge theory might profitably begin with structured vector bundles and the tensors they carry, rather than with groups. Second, a version of Ockham’s razor: if the geometry already determines the symmetry, the symmetry is doing no independent work.
@@ Appendix A Principal and associated fibre bundles
Given a vector bundle $(E,M,V)$ a covariant derivative $\nabla$ is an operator:
$\nabla:\Gamma(E)\rightarrow\Gamma(T^{*}M\otimes E)$ (A.7)
@@
$\nabla(f\kappa)={\mathrm{d}}f\otimes\kappa+f\nabla\kappa$ (A.8)
@@
Let $\nabla$ be a covariant derivative on $(E,M,V)$, $\mathbf{v}\in E_{x}$ and $\gamma(t)$ a curve in $M$ such that $\gamma(0)=x$. Then we define the parallel transport along $\gamma$ as the unique section $\mathbf{v}_{h}(t)$ of $E|_{\gamma}$ such that:
$\nabla_{\gamma^{\prime}{}}\mathbf{v}_{h}=0.$ (A.9)
The existence and uniqueness of this map is guaranteed for $\gamma\subset U$ some open subset of $M$, and it follows from properties of solutions of ordinary differential equations (cf. \@BBOPcitep\@BAP\@BBN(Kobayashi and Nomizu, 1963, Ch. II.2)\@BBCP).
Here $\nabla$ is an operator, not a tensor. But by introducing a coordinate frame or basis, we can represent it as such. This is the same as for spacetime covariant derivatives, $\nabla$: it is only upon the introduction of a frame or basis that we find an explicit representation.
To define $\Omega$ in terms of $\nabla$, we proceed in the usual way:
@@
Given a covariant derivative $\nabla$ on a vector bundle $E$, the curvature tensor is the unique multilinear bundle map
@@
$\Omega(X,Y)\kappa=\left(\nabla_{X}\nabla_{Y}-\nabla_{Y}\nabla_{X}-\nabla_{[X,Y]}\right)\kappa,$
@@
We can see the curvature then as an element of $\Omega:TM\otimes TM\otimes\mathsf{End}(E)$, i.e. as a map valued on the endomorphisms of $E$ (the fibre-linear transformations that are not necessarily automorphisms).
The trace operation is defined as $\mathsf{Tr}:\mathsf{End}(E)\rightarrow C^{\infty}(M)$, and so can be included in a Lagrangian specifying the dynamics of $\nabla$. Since $\mathsf{End}(E)$ is closed under composition, we can obtain a Lagrangian 4-form for the action:
@@
It will prove useful to know that, given any vector bundle $(E,M,V)$ the bundle of frames for $E$, called $L(E)$, is itself a principal fibre bundle $(L(E),M,GL(V))$: here elements of $\pi^{-1}(x)$ are linear frames of $E_{x}$, and $G\simeq GL(V)$ acts via $\rho$ on the typical fibres. By construction, $E\simeq L(E)\times_{\rho}V$. Now, for $G^{\prime}{}\subset G\simeq GL(V)$ we can partition the points of each orbit in $P$, $\mathcal{O}_{p}:=Gp$, into orbits of $G^{\prime}{}$. Each such choice gives a principal bundle with group $G^{\prime}{}$ and it induces further structure on the associated vector bundle, e.g. an inner product, by selecting which frames are considered orthonormal. This is also a principal fibre bundle, $(L^{\prime}{}(E),M,G^{\prime}{})$, whose structure group is a proper subgroup of the general linear group, $G^{\prime}{}\subset GL(V)$, taken to be the group that preserves the structure of $V$. This is called the bundle of admissible frames.
Sourced Yang–Mills equations and slack gauge directions
From the principal-bundle point of view, matter couples to a connection $A$ only through the infinitesimal action of the gauge Lie algebra on the matter fibres. Let $\rho:G\to U(V)$ be the unitary representation used to form the associated bundle $E=P\times_{\rho}V$, and write $\rho_{*}:\mathfrak{g}\to\mathfrak{u}(V)$ for its differential. For a Higgs field $\Phi\in\Gamma(E)$, variation of the Yang–Mills–Higgs action with respect to $A$ yields the sourced equation
$d_{A}^{*}F_{A}\;=\;J_{H}(A,\Phi),$ (A.11)
where the Higgs current $J_{H}(A,\Phi)\in\Omega^{1}(M,\mathrm{Ad}(P))$ is characterised by the Riesz identity
$\langle\alpha,J_{H}(A,\Phi)\rangle_{\mathrm{Ad}(P)}\;=\;2\,\mathrm{Re}\,\langle d_{A}\Phi,\alpha\cdot\Phi\rangle_{E}\qquad\text{for all }\alpha\in\Omega^{1}(M,\mathrm{Ad}(P)).$ (A.12)
Here $\alpha\cdot\Phi$ denotes the pointwise action induced by $\rho_{*}$: locally, if $\alpha=\alpha_{\mu}\,dx^{\mu}$ with $\alpha_{\mu}(x)\in\mathfrak{g}$, then $(\alpha\cdot\Phi)_{\mu}=\rho_{*}(\alpha_{\mu})\Phi$.
This makes explicit what happens in slack situations. If $\ker(\rho_{*})\neq 0$, then $\xi\cdot\Phi=0$ for all $\xi\in\ker(\rho_{*})$, and hence the right-hand side of (A.12) vanishes whenever $\alpha$ takes values in $\ker(\rho_{*})$. Equivalently, the current takes values only in the Lie-subalgebra that actually acts on matter:
$J_{H}(A,\Phi)\;\in\;\Omega^{1}\!\big(M,\ker(\rho_{*})^{\perp}\big)\;\subset\;\Omega^{1}(M,\mathrm{Ad}(P)),$ (A.13)
where $\ker(\rho_{*})^{\perp}$ is the orthogonal complement in $\mathfrak{g}$ with respect to the chosen Ad-invariant inner product. Accordingly, (A.11) splits into a sourced equation along $\ker(\rho_{*})^{\perp}$ and a vacuum equation along $\ker(\rho_{*})$:
$\pi_{\ker(\rho_{*})^{\perp}}(d_{A}^{*}F_{A})=J_{H}(A,\Phi),\qquad\pi_{\ker(\rho_{*})}(d_{A}^{*}F_{A})=0.$ (A.14)
Any “extra” gauge directions that do not act on the chosen matter content are automatically unsourced. They are empirically idle unless further charged fields are added.
Conversely, if the slack is only discrete—i.e. $\ker(\rho)$ is nontrivial but $\ker(\rho_{*})=0$—then the local Euler–Lagrange equation (A.11) is unchanged; the difference concerns only global issues (which principal bundles and large gauge transformations one admits).